OpenStudy (anonymous):

Scientists have discovered a solar system at a distance of 2000 AU from Earth. What is the distance, in kilometers, between the new solar system and Earth? a. 300 million kilometers b. 300 billion kilometers c. 20 million kilometers d. 20 billion kilometers

OpenStudy (anonymous):

wouldn't this be in science i mean its just a hunch

OpenStudy (anonymous):

yea it is science but nobody's responding because it has to deal w/ math ...

OpenStudy (anonymous):

ohh ummmm @mathslover

OpenStudy (anonymous):

@johnweldon1993 can you help her idk how to do this

OpenStudy (johnweldon1993):

Yeah sure :) So an astronomical Unit can be converted to KM by using the fact that \[\large \frac{1 AU}{149,597,871 km}\] So if that is true...we have 2000 AU we have \[\large \frac{1 AU}{149,597,871 km} = \frac{2000}{x}\] \[\large x = 2000 \times 149,597,871 = \space ?\]

OpenStudy (anonymous):

299,195,745,000 ? So it would be B im guessing ? :o

OpenStudy (johnweldon1993):

Indeed ^ :)

OpenStudy (anonymous):

thank you :)

OpenStudy (johnweldon1993):

Anytime :)